Reports 1-1 of 1 Clear search Modify search
AdV-INJ (ITF Mode Matching Telescope)
chiummo - 10:13 Wednesday 18 May 2022 (55889) Print this report
Comment to Matching of north arm -- estimation of north arm CP focal length (55886)

As discussed at last commissioning meeting,  the final tuning of the matching of the input beam to the arms will require the setting of a constant thermally induced lens in the compensation plates. Meanwhile, a check of the matching was performed with no CO2 laser actions.

Started from the situation as left yesterday night, with CH/DAS off. First of all a scan of the arms with green lasers was performed in order to assess the situation.

The matching of the arms looked already pretty good (see fig 1: TEM2 yesterday night, fig 2: TEM2 this morning before changing anything). The meniscus lens on SIB1 was moved by 20000 steps forward which worsen the TEM2(fig 3).

Then the position of the meniscus lens was reversed to explore the other side: moved by 40000 steps backward (fig 4). This improved again the matching comparable to initial value (around 1.5%).


To be continued after TCS actions.

Attached a text file with GPS for more accurate evaluation of the mode mismatch, the summary of the actions is reported in fig 5.





Images attached to this comment
Non-image files attached to this comment
Search Help